A 0.864 −g sample of an unknown acid requires 32.7 mL of a 0.183 M barium hydroxide solution for neutralization.

Assuming the acid is diprotic, calculate the molar mass of the acid.
Answer #1

Balanced chemical equation is:

Ba(OH)2 + H2A ---> BaA2 + 2 H2O

lets calculate the mol of Ba(OH)2

volume , V = 32.7 mL

= 3.27*10^-2 L

use:

number of mol,

n = Molarity * Volume

= 0.183*3.27*10^-2

= 5.984*10^-3 mol

According to balanced equation

mol of H2A reacted = (1/1)* moles of Ba(OH)2

= (1/1)*5.984*10^-3

= 5.984*10^-3 mol

This is number of moles of H2A

mass(H2A)= 0.864 g

use:

number of mol = mass / molar mass

5.984*10^-3 mol = (0.864 g)/molar mass

molar mass = 1.444*10^2 g/mol

Answer: 144 g/mol